OpenStudy (safia21):

KE 5 - = - and KI = 12 calculate FI FA 3

OpenStudy (safia21):

thats FL

OpenStudy (anonymous):

Is there a diagram for the question ?

OpenStudy (safia21):

nope

OpenStudy (anonymous):

any other info ?

OpenStudy (safia21):

nope!

OpenStudy (anonymous):

Geometry question ?

OpenStudy (safia21):

yea

OpenStudy (anonymous):

Not enough info. There should be a diagram, or more info provided to solve this.

OpenStudy (safia21):

there is the diagram sorry about the confusion

OpenStudy (anonymous):

FL/12 = 3/5 FL=3*12/5=36/5=7.2